Saturday, 27 September 2014

Synth Repairs

Analogue Systems RS-40 Noise, Sample and Hold, Osc

This came in as 'defective'. After a while it became obvious the noise output was not working. This was traced to an open-circuit LEVEL contol pot (10k lin) which was replaced. Still it didn't work as a sample and hold - i.e. I was expecting to see a stepped output from S\H OUT. Not having used the RS-40 before it took me a while to realise that the noise and clock signals are not normalled to the Sample and Hold circuit like the front panel markings suggest; the lines linking NOISE OUT to EXT SRC IN and INT CK OUT to EXT CK IN suggest to me these are linked in some way but no. 
After looking at the jacks it was clear that these are not normalled (I'm not sure if this is the case with all RS-40s but this one looked clean and unmodified so I'm assuming it was factory standard). I used external patch leads and to link the noise and clock outs and ins and hey presto it worked. The first thing that I saw on the scope was at each sample point the sampled signal is present for what I guess is the duration of the sample pulse (about 0.2ms). When using the RS-40 to modulate a VCO I could not tell that the noise spikes were a problem, in fact it is probably above hearing range (at least mine). 
Noise spikes

However I didn't much like it and I didn't much like not having the noise and clock not normalled to the S/H circuit. I spoke to the owner and he agreed to have a couple of simple mods applied.

The noise and clock signals are present on the jack board close to where they would have to go for normalled connections. The switched contacts were connected to ground on both EXT SRC IN and EXT CK IN jacks by solder bridges. I removed the solder bridges and fitted links as can be seen here. This worked a treat allowing it to work without patching but still allowing external signals to be patched in if need be.

Solder removed from red arrows and links applied at yellow arrows
The noise spikes 'problem' although not audible was fixed by placing a 10nF cap across the S/H output amplifier's feedback resistor (27k) making it a low-pass filter with a cutoff of 590Hz. This got rid of the spikes without adding noticeable slew.

10nF capacitor

Spike free steps

No comments:

Post a Comment